A horizontal force of 10 N is applied to a 4.0 kg block that is at rest on a perfectly smooth, level surface. Find the speed of the block and how far it has gone after 6.0 s?
1
Expert's answer
2020-12-06T17:46:01-0500

We have given:

F=10N, m=4.0kg, Vinitial=0, and we need to find:

Vfinal-speed of block, S-reached distance after 6 sec.

We can easily find first acceleration, then speed and distance.

F="m\times a"=10N -->a=2.5"\frac{m}{s^2}" ,

Now we find speed of block after 6 seconds:

Vfinal=Vinitial+"a\times t", Vinitial=0 because it was at rest then it moved.

Vfinal="2.5\times 6" =15"\frac{m}{s}"

S=Vinitial"\times t" + "a\times t^2" =0+90=90 m.

We find asked things.